Grade 11Mechanics

The resultant of two forces 2P and √2P is √10P. the angle betweeen the forces is
  1. 30
  2. 60
  3. 45
  4. 90

Yash
8 Years agoGrade 11

1 Answer

Khimraj
ApprovedApproved Tutor Answer8 Years ago
resultant of two vector is given by \sqrt{A^{2} + B^{2} + 2ABcos\Theta }
where \Theta is angle between the vectors.
so
\sqrt{10}P = \sqrt{4 + 2 + 4\sqrt{2}cos\Theta }P
squaring both side
10 P2 = (6 + 4\sqrt{2}cos\Theta)P2
4 = 4\sqrt{2}cos\Theta
SO \Theta = 45
Hope it clears. If you like answer then please approve it.